OpenStudy (anonymous):

9th Grade- I really need help with these questions. I have absolutely no idea on the last one but the first to i think I got them right. 1. 2/3(c-18)= 7 2. a/20 = 4/15 = 9/15 3. b/3 = 1/8 =19 I think the first one is 28.5 and the second one 20/3 but I have no clue on the third. If you can, please show the work and steps! Thanks!!!

OpenStudy (across):

Hint:\[\begin{align}\frac23(c-18)&=7\\2(c-18)&=21\\c-18&=\frac{21}2\\c&=\frac{21}2+18=\frac{21+36}2=28.5\end{align}\]You're correct. Keep solving them the way you're doing so right now.

OpenStudy (across):

The third one makes no sense; you're claiming that\[\frac18=19.\]This is false.

OpenStudy (across):

Did you mean\[\frac b3+\frac18=19?\]

OpenStudy (anonymous):

Sorry, #3 is b/3 + 1/8 =19

OpenStudy (across):

Well, you have that\[\frac b3+\frac18=19.\]To isolate \(b\), you move \(1/8\) to the other side:\[\frac b3=19-\frac18=\frac{152-1}{8}=\frac{151}{8}.\]Then you multiply both sides by \(3\):\[b=\frac{151\cdot3}{8}=\frac{453}8.\]

OpenStudy (anonymous):

Thank you SO much! Math is NOT my strong subject you really helped a lot!!!

OpenStudy (across):

You're always welcome.